OpenStudy (anonymous):

Solve for x (x-2)^-4=625

OpenStudy (anonymous):

Do you know of x^-2 = 1/(x^2)?

OpenStudy (anonymous):

no I think the answer is x=7 is that correct

geerky42 (geerky42):

\((x-2)^4 =\dfrac{1}{625}\) \(\sqrt[4]{(x-2)^4} = \sqrt[4]{\dfrac{1}{625}}\) \(|x-2| = \dfrac{1}{\sqrt[4]{625}} = \dfrac{1}{5}\) \(|x-2| = \dfrac{1}{5}\)

geerky42 (geerky42):

There are two answers.

geerky42 (geerky42):

x= 7 is not correct.

OpenStudy (anonymous):

Ok I have math processong errors when you gave me thAt first reply and how do you get two solutions

geerky42 (geerky42):

Look at last step. There is absolute value.

geerky42 (geerky42):

Do you know how to get rid of absolute value?

OpenStudy (anonymous):

I can't see your last step because it all says math processing error

OpenStudy (anonymous):

So then does x=.5 and x=-2 is that what you are meaning what the answers are

OpenStudy (anonymous):

Is that what you are meaing

OpenStudy (anonymous):

Can someone tell me if x=-2 and x=.2 are the answers to this problem